# FineScope: SAE-guided Data Selection Enables Domain-Specific LLM Pruning & Fine-Tuning

### v2

@@ Abstract
Large Language Models (LLMs) are typically trained on diverse, general-purpose datasets, enabling broad generalization but incurring substantial computational costs. However, real-world applications often require efficient models tailored to narrow, domain-specific tasks. In such settings, large model capacity and generality are unnecessary, and traditional fine-tuning pipelines struggle under resource constraints. We introduce FineScope, a framework that addresses this challenge by tightly coupling domain-aware data selection with model pruning and fine-tuning. Starting from a small set of user-provided seed examples, FineScope trains sparse autoencoders (SAEs) on intermediate model activations to automatically extract semantically aligned examples from large unlabeled corpora. The curated dataset then guides structured pruning to preserve domain-relevant substructures and supports self-distillation fine-tuning to recover task-specific performance. Experiments across STEM, humanities, social sciences, math, and coding domains show that FineScope consistently outperforms baseline fine-tuning approaches while enabling up to $35\%$ parameter pruning. On math reasoning tasks, it achieves an average improvement of  11.50 points across pruned models.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated strong generalization across a wide range of tasks due to their training on broad and diverse datasets(borealis2024highlevel, ; b1, ). However, most real-world applications run in resource-constrained environments (upstagesolar, ; llama3, ; nvidia70, ) and require models to perform well on a narrow set of domain-specific tasks. In these cases, the broad capabilities of large models are often unnecessary, and much of their computational cost is spent on supporting functions that are irrelevant to the target domain. This creates a growing need for models that are both efficient and specialized for specific applications.
@@
Figure 1: Overall accuracy (with different model pruning ratio) on the STEM, Social Sciences, and Humanities domains after fine-tuning with Self-Instruct dataset (selfinstruct, ) and FineScope dataset. Despite its smaller size, FineScope sustains higher accuracy under aggressive pruning, underscoring its data efficiency and quality.
@@
Despite significant progress in model compression techniques such as pruning and quantization, effective domain adaptation still depends heavily on the availability of high-quality, domain-specific data. In an ideal setting, a well-constructed dataset tailored to the target domain would be used for fine-tuning (csr, ; mcq, ; problemstatementpaper, ). However, such datasets are rarely available in practice. Most existing approaches rely on general-purpose instruction datasets (alpaca, ) or manually curated corpora, which can be noisy, expensive to build, or misaligned with the intended application. As a result, models are often fine-tuned with suboptimal data which hinders performance recovery after compression.
@@
A key challenge in domain adaptation is the limited availability of high quality, domain aligned data for fine tuning. While much of the existing work focuses on model compression and adaptation techniques, these approaches often assume access to suitable data or treat data selection as a secondary concern. In practice, the effectiveness of compressed models depends heavily on the quality and relevance of the fine tuning dataset (lima, ). Without data that closely reflects the target domain, even the most advanced compression strategies are unlikely to maintain strong performance.
@@
To address this challenge, we propose that data selection should be treated as a central part of the adaptation process. Inspired by findings that data quality can be more impactful than quantity (lima, ), we argue that a small, carefully chosen subset of relevant data can support strong performance, even in heavily compressed models. Rather than treating data curation and model optimization as separate steps, we design an approach that connects them, allowing the data to guide both pruning and fine-tuning within a unified framework.
@@
We present FineScope, a framework that automates domain-specific data selection and integrates it with model pruning and fine-tuning. FineScope begins with a small set of user-provided seed examples and uses a Sparse Autoencoder (SAE), trained on intermediate activations of a pretrained LLM, to identify semantically relevant samples (anthropy, ; kissane2024interpreting, ; yan2024encourage, ) from a large unlabeled corpus. These curated samples form a compact, high-quality dataset that reflects the target domain and is used to guide structured pruning. A modified self-distillation fine-tuning step then helps recover any task-relevant knowledge lost during compression.
@@ 2 Related Work
(1) Domain-Specific Language Models Recent efforts have adapted large language models to specialized domains by training or fine-tuning them on domain-specific datasets. Examples include PharmaGPT (pharmagpt, ) for biomedical applications, SaulLM (saullm, ) for legal tasks, Shai (shai, ) for asset management, BloombergGPT (bloomberg, ) for financial analysis, and MedPalm (medpalm, ) for medical question answering. Additional models such as ClimateBERT (climatebert, ), ChatLaw (chatlaw, ), and FinGPT (fingpt, ) focus on areas such as climate science, legal reasoning, and financial modeling. While effective, these models typically rely on access to large, high-quality domain datasets and require full-scale retraining of billion-parameter models. Few methods address the challenges of adapting models efficiently, particularly in settings with limited compute and annotated data. In contrast, our work focuses on enabling domain adaptation by automatically selecting relevant data from large unlabeled corpora and compressing models to reduce computational requirements without sacrificing task performance.
@@
(2) Pruning Pruning is a widely used technique for reducing the size and computational cost of language models by removing parameters that have minimal impact on performance. Motivated by the lottery ticket hypothesis (lottery, ), many methods aim to identify smaller subnetworks within large models that can be trained or fine-tuned to match the original model’s accuracy. Structured pruning, in particular, removes entire architectural components such as attention heads or feedforward blocks, resulting in models that are both compact and efficient in hardware implementation.
@@
Existing pruning methods, both structured and unstructured (structured, ; taskstructured, ; other1, ; other2, ; other3, ; shearedllama, ), are typically designed to maintain general-purpose capabilities and are applied independently of the data used for downstream fine-tuning. As a result, they may not fully account for the specific needs of domain adaptation. In our approach, pruning is informed by data that is automatically selected for its relevance to a target domain. By integrating data selection into the pruning process, where we prune the model with respect to the domain-specific dataset, we aim to retain subnetworks that are more closely aligned with domain-specific behavior.
@@
(3) Neural Representation Alignment While data-centric adaptation has gained increasing attention, most methods treat the model as fixed and focus solely on selecting or generating training examples. Instruction tuning and self-training approaches (lima, ; alpaca, ; selfinstruct, ) improve performance by using curated prompts or synthetic data but do not adapt the underlying model structure. Retrieval-based methods (bricken2023decomposing, ; gadre2023datacomp, ) select data based on surface-level similarity or metadata, which may not accurately reflect the model’s internal understanding of domain relevance.
@@
Recent studies suggest that intermediate model activations encode meaningful signals related to both task characteristics and input semantics (anthropy, ). However, this insight has rarely been applied to guide either data selection or model compression. Our work builds on this idea by using a Sparse Autoencoder trained on the top-$k$ intermediate activations of a pretrained model to estimate domain relevance. This allows us to identify training data that aligns with the model’s internal representations and use it to inform both pruning and fine-tuning.
@@
Figure 2: Overview of the FineScope : (1) Dataset Curation: (a) Sparse Autoencoder (SAE) is trained on the top-$K$ activations of a pretrained LLM and then used it to extract embedding from datasets,(b)Domain-specific dataset is curated by computing cosine similarity between target domain and the samples in the larger dataset. (2) Pruning and Fine-Tuning: (c) Structured pruning is done w.r.t. selected dataset; (d) Fine-tuned the pruned model using modified self distillation. Here, $U$ denotes the larger dataset, while $D_{t}$ represents the target domain dataset. The corresponding embeddings are denoted by $e_{u}$ and $e_{t}$ respectively.
@@ 3 FineScope
3.1 Sparse-Autoencoder-Guided Domain-Specific Dataset Curation
@@ 3.1.1 Training SAE
A Sparse Autoencoder (SAE) is a neural network designed to learn compressed representations of input data while enforcing sparsity constraints on the hidden units (EleutherAI_SAE, ). In our framework, SAEs serve as a mechanism for extracting domain-relevant features from pretrained LLM activations to identify domain-relevant samples from a large corpus. Instead of operating directly on the raw model outputs, we train SAEs on activations from intermediate layers of the LLM, allowing us to capture a structured, low-dimensional representation of the underlying knowledge encoded in the model, as shown in Figure 2a. Additionally, since processing all activations are computationally infeasible for the large corpus, our SAEs are adapted to learn only the representative activations in a way that highlights the most significant neurons, improving interpretability while discarding less relevant signals.
In decoder-only transformer models such as GPT-2 and LLaMa, the activation flow can be formally defined as follows (e2e, ):
$act^{(l)}(x)=f^{(l)}\left(act^{(l-1)}(x)\right),\text{for }l=1,\dots,L-1$ (1)
@@
To train the SAE, we extract activations $act^{(l)}(x)$ from a selected layer of the pretrained LLM and feed them into the encoder network:
$\text{Enc}\left(act^{(l)}(x)\right)=\text{ReLU}\left(W_{e}act^{(l)}(x)+b_{e}\right).$ (3)
@@
$\text{SAE}\left(act^{(l)}(x)\right)=Dec^{\top}\text{Enc}\left(act^{(l)}(x)\right)+b_{dec}.$ (4)
Here, Enc represents the SAE’s encoder, $W_{e}$ and $b_{e}$ are the encoder’s weights and biases, and $Dec^{\top}$ and $b_{dec}$ correspond to the decoder parameters. The SAE is trained to minimize the reconstruction loss:
$\mathcal{L}_{\mathrm{SAE}}=\left\lVert\mathrm{SAE}\!\big(\mathrm{act}^{(l)}(x)\big)-\mathrm{act}^{(l)}(x)\right\rVert_{2}^{2}+\lambda\left\lVert\mathrm{Enc}\!\big(\mathrm{act}^{(l)}(x)\big)\right\rVert_{1}.$ (5)
where the second term enforces sparsity by penalizing the activation magnitudes of the encoder
3.1.2 Top-K Activation Selection for Efficient SAE Training
Rather than training the SAE on full-layer activations, we apply a Top-K filtering mechanism to select the most important activations before feeding them into the encoder. Given dataset $D$ containing $m$ samples, $D=\{x_{1},x_{2},\dots,x_{m}\}$, we compute the $K$ most significant activations based on the gradient magnitude:
$\text{TopK}(act^{(l)}(x),K)=\{a_{i}\in act^{(l)}(x)\mid\\ a_{i}\text{ is one of the }K\text{ largest }\left|\frac{\partial act^{(l)}_{i}(x)}{\partial x}\right|\}.$ (6)
Using only the $TopK$ activations reduces the input dimensionality of the SAE, significantly lowering computational overhead and enabling efficient training on large-scale corpora. Additionally, filtering out noisy activations improves generalization by focusing the SAE on the most informative neurons, enhancing interpretability of extracted features aligned with domain-relevant information.
The encoder then operates on these filtered activations as:
$\text{Enc}\left(\text{TopK}(act^{(l)}(x),K)\right)=\text{ReLU}\left(W_{e}\cdot\text{TopK}(act^{(l)}(x),K)+b_{e}\right).$ (7)
For each layer $l$ in the LLM, we train a separate SAE, resulting in a total of $L$ SAEs. Once trained, these SAEs serve as feature extractors for data curation.
@@ 3.1.3 Dataset Curation
To identify samples in $U$ that are most similar to $D_{t}$, we compute the cosine similarity between each embedding $e_{u}\in E_{U}$ and every embedding $e_{t}\in E_{t}$. Using this similarity measure, we select samples $x_{u}\in U$ whose embeddings fall within the Top-K highest cosine similarity scores when compared to the embeddings in $E_{t}$. Specifically, we define the selected dataset $D_{s}$ as:
$D_{s}=\{x_{u}\in U\mid e_{u}\in\\ \text{TopK}(\{\text{CosSim}(\text{SAE}(x_{u}),e_{t}):e_{t}\in E_{t}\})\}.$ (8)
where $\text{CosSim}(e_{u},e_{t})$ denotes the cosine similarity.
The final dataset $D_{s}$ contains samples from $U$ that are most semantically similar to the target domain $D_{t}$ based on the cosine similarity of their SAE embeddings. To ensure that $D_{s}$ remains a high-quality domain-specific dataset. In our evaluation, the value of $K$ across all SAEs is set to $100$ for consistency in selection.
3.2 Pruning and Fine-Tuning with Modified SDFT

### v3

# arXiv:2505.00624v3[cs.CL] 27 Feb 2026

@@ Abstract
Large Language Models (LLMs) are typically trained on diverse, general-purpose datasets, enabling broad generalization but incurring substantial computational costs. However, real-world applications often require efficient models tailored to narrow, domain-specific tasks. In such settings, large model capacity and generality are unnecessary, and traditional fine-tuning pipelines struggle under resource constraints. We introduce FineScope, a framework that addresses this challenge by tightly coupling domain-aware data selection with model pruning and fine-tuning. Starting from a small set of user-provided seed examples, FineScope trains sparse autoencoders (SAEs) on intermediate model activations to automatically extract semantically aligned examples from large unlabeled corpora. The curated dataset then guides structured pruning to preserve domain-relevant substructures and supports self-distillation fine-tuning to recover task-specific performance. Experiments across STEM, humanities, social sciences, math, and coding domains show that FineScope consistently outperforms baseline fine-tuning approaches while enabling up to $35\%$ parameter pruning. On math reasoning tasks, it achieves an average improvement of  11.50 points across pruned models. Code will be available.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated strong generalization across a wide range of tasks due to their training on broad and diverse datasets(RBC Borealis AI, 2024; Adler et al., 2024). However, most real-world applications run in resource-constrained environments (Kim et al., 2023; Dubey et al., 2024; Wang et al., 2024) and require models to perform well on a narrow set of domain-specific tasks. In these cases, the broad capabilities of large models are often unnecessary, and much of their computational cost is spent on supporting functions that are irrelevant to the target domain. This creates a growing need for models that are both efficient and specialized for specific applications.
@@
Figure 1: Overall accuracy (with different model pruning ratio) on the STEM, Social Sciences, and Humanities domains after fine-tuning with Self-Instruct dataset (Wang et al., 2022) and FineScope dataset. Despite its smaller size, FineScope sustains higher accuracy under aggressive pruning, underscoring its data efficiency and quality.
@@
Despite significant progress in model compression techniques such as pruning and quantization, effective domain adaptation still depends heavily on the availability of high-quality, domain-specific data. In an ideal setting, a well-constructed dataset tailored to the target domain would be used for fine-tuning (Talmor et al., 2018; Clark et al., 2018; Gu et al., 2024). However, such datasets are rarely available in practice. Most existing approaches rely on general-purpose instruction datasets (Taori et al., 2023) or manually curated corpora, which can be noisy, expensive to build, or misaligned with the intended application. As a result, models are often fine-tuned with suboptimal data which hinders performance recovery after compression.
@@
A key challenge in domain adaptation is the limited availability of high quality, domain-aligned data for fine tuning. While much of the existing work focuses on model compression and adaptation techniques, these approaches often assume access to suitable data. In practice, the effectiveness of compressed models depends heavily on the quality and relevance of the fine tuning dataset (Zhou et al., 2024). Without data that closely reflects the target domain, even the most advanced compression strategies are unlikely to maintain strong performance.
@@
To address this challenge, we propose that data selection should be treated as a central part of the adaptation process. Inspired by findings that data quality can be more impactful than quantity (Zhou et al., 2024), we argue that a small, carefully chosen subset of relevant data can support strong performance, even in heavily compressed models. Rather than treating data curation and model optimization as separate steps, we design an approach that connects them, allowing the data to guide both pruning and fine-tuning within a unified framework.
@@
We present FineScope, a framework that automates domain-specific data selection and integrates it with model pruning and fine-tuning. FineScope begins with a small set of user-provided seed examples and uses a Sparse Autoencoder (SAE), trained on intermediate activations of a pretrained LLM, to identify semantically relevant samples (Templeton et al., 2024; Kissane et al., 2024; Yan et al., 2024) from a large unlabeled corpus. These curated samples form a compact, high-quality dataset that reflects the target domain and is used to guide structured pruning. A modified self-distillation fine-tuning step then helps recover any task-relevant knowledge lost during compression.
@@ 2 Related Work
(1) Domain-Specific Language Models Recent efforts have adapted large language models to specialized domains by training or fine-tuning them on domain-specific datasets. Examples include PharmaGPT (Chen et al., 2024) for biomedical applications, SaulLM (Colombo et al., 2024) for legal tasks, Shai (Guo et al., 2023) for asset management, BloombergGPT (Wu et al., 2023) for financial analysis, and MedPalm (Singhal et al., 2023) for medical question answering. Additional models such as ClimateBERT (Webersinke et al., 2021), ChatLaw (Cui et al., 2023), and FinGPT (Yang et al., 2023) focus on areas such as climate science, legal reasoning, and financial modeling. While effective, these models typically rely on access to large, high-quality domain datasets and require full-scale retraining of billion-parameter models. Few methods address the challenges of adapting models efficiently, particularly in settings with limited compute and annotated data. In contrast, our work focuses on enabling domain adaptation by automatically selecting relevant data from large unlabeled corpora and compressing models to reduce computational requirements without sacrificing task performance.
@@
(2) Pruning Pruning is a widely used technique for reducing the size and computational cost of language models by removing parameters that have minimal impact on performance. Motivated by the lottery ticket hypothesis (Frankle and Carbin, 2018), many methods aim to identify smaller subnetworks within large models that can be trained or fine-tuned to match the original model’s accuracy. Structured pruning, in particular, removes entire architectural components such as attention heads or feedforward blocks, resulting in models that are both compact and efficient in hardware implementation.
@@
Existing pruning methods, both structured and unstructured (Wang et al., 2019; Xia et al., 2022; Zafrir et al., 2021; Kurtic et al., 2022; Ashkboos et al., 2024; Xia et al., 2023), are typically designed to maintain general-purpose capabilities and are applied independently of the data used for downstream fine-tuning. As a result, they may not fully account for the specific needs of domain adaptation. In our approach, pruning is informed by data that is automatically selected for its relevance to a target domain. By integrating data selection into the pruning process, where we prune the model with respect to the domain-specific dataset, we aim to retain subnetworks that are more closely aligned with domain-specific behavior.
@@
(3) Neural Representation Alignment While data-centric adaptation has gained increasing attention, most methods treat the model as fixed and focus solely on selecting or generating training examples. Instruction tuning and self-training approaches (Zhou et al., 2024; Taori et al., 2023; Wang et al., 2022) improve performance by using curated prompts or synthetic data but do not adapt the underlying model structure. Retrieval-based methods (Bricken et al., 2023; Gadre et al., 2023) select data based on surface-level similarity or metadata, which may not accurately reflect the model’s internal understanding of domain relevance.
@@
Recent studies suggest that intermediate model activations encode meaningful signals related to both task characteristics and input semantics (Templeton et al., 2024). However, this insight has rarely been applied to guide either data selection or model compression. Our work builds on this idea by using a Sparse Autoencoder trained on the top-$k$ intermediate activations of a pretrained model to estimate domain relevance. This allows us to identify training data that aligns with the model’s internal representations and use it to inform both pruning and fine-tuning.
@@
Figure 2: Overview of the FineScope: (1) Dataset Curation: (a) Sparse Autoencoder (SAE) is trained on the top-$K$ activations and then used it to extract embedding from datasets, (b) Domain-specific dataset is curated by computing cosine similarity between target domain and the samples in the larger dataset. (2) Pruning and Fine-Tuning: (c) Structured pruning is done w.r.t. selected dataset; (d) Fine-tuned the pruned model using modified self distillation. Here, $U$ denotes the larger dataset, while $D_{t}$ represents the target domain dataset. The corresponding embeddings are denoted by $e_{u}$ and $e_{t}$ respectively.
@@ 3 FineScope
3.1 Methods
@@ 3.1.1 Training SAE
A Sparse Autoencoder (SAE) is a neural network designed to learn compressed representations of input data while enforcing sparsity constraints on the hidden units (EleutherAI, 2024). In our framework, SAEs serve as a mechanism for extracting domain-relevant features from pretrained LLM activations to identify domain-relevant samples from a large corpus. (We discuss why we used SAE encoder code as the embedding for retrieval in the supplement; see Section 7.1.)
Instead of operating directly on the raw model outputs, we train SAEs on activations from intermediate layers of the LLM, allowing us to capture a structured, low-dimensional representation of the underlying knowledge encoded in the model, as shown in Figure 2a. Additionally, since processing all activations is computationally infeasible for a large corpus, our SAEs are adapted to learn only the representative activations in a way that highlights the most significant neurons, improving interpretability while discarding less relevant signals.
In decoder-only transformer models such as GPT-2 and LLaMa, the activation flow can be formally defined as follows (Braun et al., 2024):
$act^{(l)}(x)=f^{(l)}\left(act^{(l-1)}(x)\right),\text{for }l=1,\dots,L-1.$ (1)
@@
To train the SAE, we extract activations $act^{(\ell)}(x)$ from a selected layer of the pretrained LLM and feed them into the encoder network:
$\mathrm{Enc}\!\left(act^{(\ell)}(x)\right)=\mathrm{ReLU}\!\left(W_{e}\,act^{(\ell)}(x)+b_{e}\right).$ (3)
@@
$\mathrm{SAE}\!\left(act^{(\ell)}(x)\right)=\mathrm{Dec}^{\top}\mathrm{Enc}\!\left(act^{(\ell)}(x)\right)+b_{\mathrm{dec}}.$ (4)
We train the SAE with a reconstruction loss plus an $\ell_{1}$ sparsity penalty on encoder activations:
$\mathcal{L}_{\mathrm{SAE}}=\left\lVert\mathrm{SAE}\!\big(act^{(\ell)}(x)\big)-act^{(\ell)}(x)\right\rVert_{2}^{2}+\lambda\left\lVert\mathrm{Enc}\!\big(act^{(\ell)}(x)\big)\right\rVert_{1}.$ (5)
The second term enforces sparsity by penalizing the activation magnitudes of the encoder. In our implementation, we use a scale-normalized reconstruction objective and additional stabilization terms; see the Supplementary Material for full loss details (Section 6).
3.1.2 Top-$K$ Activation Selection for SAE Training
Training an SAE on the full activation width $d_{\ell}$ is costly. We therefore preselect, for each layer $\ell$, a fixed subset of $K_{\text{act}}\ll d_{\ell}$ activation coordinates and train the SAE only on those coordinates.
Let $x$ be an input sequence, and let $e(x)\in\mathbb{R}^{T\times d_{\text{model}}}$ denote the embedding activations (treated as a differentiable input tensor; the embedding parameters are frozen). Let $a^{(\ell)}(x)\in\mathbb{R}^{d_{\ell}}$ be a sequence-level activation vector at a chosen hookpoint in layer $\ell$ . We define the per-coordinate input-sensitivity as
$s^{(\ell)}_{j}(x)\;=\;\left\|\frac{\partial a^{(\ell)}_{j}(x)}{\partial e(x)}\right\|_{F},$ (6)
which corresponds to the Frobenius norm of the Jacobian row associated with coordinate $j$. In the supplement, we provide additional intuition for Jacobian-sensitivity as a coordinate-selection criterion (Section 7.2), and an ablation comparing alternative Top-K selectors (Section 8).
Efficient estimation via randomized JVPs.
Computing Eq. (6) exactly for all $j$ would require access to the full Jacobian. Instead, we estimate the squared saliency using a Hutchinson-style estimator that only requires $R$ Jacobian vector products (JVPs), with small $R$ (e.g., $R\in\{1,2,4\}$), independent of $d_{\ell}$.
Let $J^{(\ell)}(x)=\frac{\partial a^{(\ell)}(x)}{\partial\mathrm{vec}(e(x))}\in\mathbb{R}^{d_{\ell}\times(Td_{\text{model}})}$, and sample probe vectors $r^{(p)}\in\{\pm 1\}^{Td_{\text{model}}}$ with i.i.d. Rademacher entries. For each probe, we compute a JVP
$u^{(p)}(x)\;=\;J^{(\ell)}(x)\,r^{(p)},$ (7)
which can be obtained by forward-mode automatic differentiation. Then $\mathbb{E}\big[(u^{(p)}_{j}(x))^{2}\big]=\left\|J^{(\ell)}_{j,:}(x)\right\|_{2}^{2}$, so we form the estimator
$\widehat{s}^{(\ell)}_{j}(x)^{2}\;=\;\frac{1}{R}\sum_{p=1}^{R}\left(u^{(p)}_{j}(x)\right)^{2}.$ (8)
Global coordinate ranking.
We aggregate over a reference corpus $D_{0}$:
$W^{(\ell)}_{j}\;=\;\mathbb{E}_{x\sim D_{0}}\big[\widehat{s}^{(\ell)}_{j}(x)\big]\;\approx\;\frac{1}{|D_{0}|}\sum_{x\in D_{0}}\widehat{s}^{(\ell)}_{j}(x),$ (9)
and select a fixed index set
$I^{(\ell)}=\mathrm{TopK}\big(\{W^{(\ell)}_{j}\}_{j=1}^{d_{\ell}},K_{\text{act}}\big).$ (10)
During SAE training we restrict activations to these coordinates:
$\tilde{a}^{(\ell)}(x)=a^{(\ell)}(x)\big[I^{(\ell)}\big]\in\mathbb{R}^{K_{\text{act}}},$ (11)
and train a layer-specific SAE on $\tilde{a}^{(\ell)}(x)$.
@@ 3.1.3 Dataset Curation
To identify samples in $U$ that are most similar to $D_{t}$, we first compute SAE embeddings $e_{u}=\mathrm{SAE}(x_{u})$ for $x_{u}\in U$ and $e_{t}=\mathrm{SAE}(x_{t})$ for $x_{t}\in D_{t}$. For each sample $x_{u}\in U$, we define its similarity to the target domain as
$s(x_{u})=\max_{x_{t}\in D_{t}}\mathrm{CosSim}\!\left(\mathrm{SAE}(x_{u}),\mathrm{SAE}(x_{t})\right).$ (12)
We then construct the curated dataset by selecting the Top-$M$ samples in $U$ with the highest similarity scores:
$D_{s}=\operatorname{TopM}_{x_{u}\in U}\big(s(x_{u}),M\big).$ (13)
The final dataset $D_{s}$ contains samples from $U$ that are most semantically similar to the target domain $D_{t}$ based on the cosine similarity of their SAE embeddings. To ensure that $D_{s}$ remains a high-quality domain-specific dataset. In our evaluation, the value of $M$ across all SAEs is set to $100$ for consistency in selection.
3.2 Pruning and Fine-Tuning